Solve for R: V=12R/(r+R)
Answer by rothauserc(4718)

V = 12R/(r+R)
:
V(r+R) = 12R
:
Note I assume r+R is not = 0
:
Vr +VR = 12R
:
12R -VR = Vr
:
R(12-V) = Vr
:
R = Vr/(12-V)
: